OpenStudy (anonymous):

Medals! Use the Laws of Exponents to simplify the expression four to the power of forty over four to the power of twenty Explain how you found your answer and write the answer as an exponent.

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

also ill fan if you can do this one. Use the Laws of Exponents to simplify the expression four to the power of forty over four to the power of twenty Explain how you found your answer and write the answer as an exponent.

OpenStudy (anonymous):

@bibby

OpenStudy (campbell_st):

when you have the same base subtract the powers... \[\frac{x^a}{x^b} = x^{a - b}\] use this law for the answer

OpenStudy (anonymous):

English?!?!?!?! @campbell_st

OpenStudy (anonymous):

do u still need help on this honey?

OpenStudy (shamil98):

\[\huge \frac{ 4^{40} }{ 4^{20} } = 4^{40-20} = 4^{20}\]